# EFX Exists for Three Types of Agents

### v1

@@ 3 Three sources in $E_{X}$
As the leading agents hold the minimum valued bundle in their respective groups, they envy all the agents from their own group and are never envied by anyone in their own group. We consider the case when no leading agent is envied by anyone. That is, $a_{1},b_{1}$ and $c_{1}$ are the sources in $E_{X}$.
@@
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If the set of leading agents $L=\{a_{1},b_{1},c_{1}\}$ are sources in the envy graph $E_{X}$, then there exists an allocation $Y$ such that $Y\succ X$. Furthermore, $Y(L)\succ X(L)$ and $Y(\mathcal{N}\setminus{L})=X(\mathcal{N}\setminus{L})$.
@@
In this case, we only update the bundles of the leading agents, while the remaining agents retain their old bundles. After the update, each leading agent prefers their new bundle over their old ones. Additionally, the leading agents do not envy the new bundles of other leading agents, even when compared to their own old bundles. Thus, the non-leading agents in a group (who keep their old bundles) do not strongly envy any agent in other groups. As a result, the only strong envies that remain are within the groups, which are resolved using Lemma 4. Therefore, the new allocation we obtain is an EFX allocation. Furthermore, the new bundles satisfy $Y(L)\succ X(L)$ and $Y(\mathcal{N}\setminus L)=X(\mathcal{N}\setminus L)$.
Proof.
From Proposition 4, we know that the $g$-champion graph $M_{X}$ must contain a cycle of length at most three, involving agents $a_{1},b_{1}$, and $c_{1}$. If any of the leading agents self $g$-champions, using Corollary 1, we get a Pareto-dominating EFX allocation $Y$. If the $g$-champion cycle is of length $2$, then by applying Lemma 5, we get our desired allocation. Therefore, assume that $M_{X}$ has a cycle of length $3$. Without loss of generality, assume that $(a_{1},c_{1})$,$(c_{1},b_{1})$ and $(b_{1},a_{1})$ are $g$-champion edges in $M_{X}$ as shown in Figure 2. Since these agents do not envy each other, the $g$-champion edges break the bundle into top and bottom halves as shown below.
@@ 1. Case 1
    We allocate $T_{c}\cup B_{b}$ to agent $a_{1}$, $T_{a}\cup B_{c}$ to agent $b_{1}$, and $T_{b}\cup B_{a}$ to agent $c_{1}$. These bundles correspond to the most preferred top and bottom halves for each agent (refer to Conditions 4, 5, and 6). Therefore, all three agents strictly improve. Furthermore, according to agent $a_{1}$, we have $T_{c}\cup B_{b}>_{a}T_{a}\cup B_{a}=X_{a_{1}}>_{a}T_{a}\cup B_{c}=Y_{b_{1}}$. That is, $a_{1}$ does not envy $b_{1}$ even with respect to her old bundle $X_{a_{1}}$. Therefore, no agent in $\mathbb{A}$ envy $b_{1}$. Similarly, we have $X_{a_{1}}>_{a}T_{b}\cup B_{a}=Y_{b_{1}}$. That is, no agent from $\mathbb{A}$ envies $b_{1}$. Similarly we can show that no agents from either $\mathbb{B}$ or $C$ envy any of $a_{1},b_{1}$ or $c_{1}$. If any strong envy remains, it must be confined to agents within their respective groups. Hence, we apply Lemma 4 to resolve these envies.
@@ 2. Case 2
    None of the three predicates holds: Thus $B_{a}>_{a}B_{b}$, $B_{b}>_{b}B_{c}$, and $B_{c}>_{c}B_{a}$. We then cyclically shift the top halves of these three bundles, allocating $T_{c}\cup B_{a}$ to agent $a_{1}$, $T_{a}\cup B_{b}$ to agent $b_{1}$, and $T_{b}\cup B_{c}$ to agent $c_{1}$. Since $T_{c}>_{a}T_{a}$ and $T_{c}>_{a}T_{b}$, agent $a_{1}$ strictly improves and does not envy either $b_{1}$ or $c_{1}$ as they received less preferred top halves according to $a_{1}$. Similarly, both $b_{1}$ and $c_{1}$ also experience improvements and do not envy any agents from $\{a_{1},b_{1},c_{1}\}$. Furthermore, by applying the inequalities 1,2, and 3 we have $T_{b}>_{a}T_{a}$ which implies $T_{b}\cup B_{b}>_{a}T_{a}\cup B_{b}=Y_{b_{1}}$. Along with the fact that $a_{1}$ did not previously envy $b_{1}$, we have $X_{a_{1}}>_{a}X_{b_{1}}>_{a}Y_{b_{1}}$. Therefore agent $a_{1}$ does not envy $b_{1}$ even with respect to her old bundle $X_{a_{1}}$. Therefore, no agent in $\mathbb{A}$ envy $b_{1}$. In an analogous way we can show that none of the agents envy any of $a_{1},b_{1}$ or $c_{1}$. If any strong envy remains, it must be confined to agents within their respective groups. Hence, we apply Lemma 4 to resolve these envies.
@@ 4. Case 4
        $\mathbf{S>_{c}X_{c_{1}}}$: Then, we allocate the bundles as follows. Agent $a_{1}$ gets $T_{c}\cup B_{a}$, $b_{1}$ retains $T_{b}\cup B_{b}$ and $c_{1}$ gets $S$. From Condition 4 we have $T_{c}\cup B_{a}>_{a}T_{a}\cup\mathbb{B}_{a}$ and from this case assumtion we know that $S>_{c}X_{c_{1}}$. Therefore agents $a_{1}$ and $c_{1}$ receive strict improvements while agent $b_{1}$ retains her old bundle.
@@
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If the set of leading agents $L=\{a_{1},b_{1},c_{1}\}$ are all the sources in the envy graph $E_{X}$, then there exists an EFX allocation $Y$, possibly partial, such that at least one of the following holds:
@@
If, however, $b_{1}$ $g$-champions $a_{1}$, then $\langle(a_{1},b_{1}),(b_{1},a_{1})\rangle$ is a $\phi$-improving exchange, thus we are done. The only remaining possibility is that $c_{1}$ $g$-champions $a_{1}$. We now consider cases based on who $g$-champions $c_{1}$.
@@ 3. Case 3
      Agent $c_{1}$ envies some $a_{i}\in\mathbb{A}\setminus\{a_{1}\}$: Then, we exchange the bundles of agents $a_{i}$ and $c_{1}$. Note that there may not be envy from $a_{i}$ to $c_{1}$, But since $a_{1}$ envies $c_{1}$, the bundle of $c_{1}$ is better than bundle of $a_{1}$ according to $v_{a}$. Now, all the strong envies are within the groups. Therefore, we apply Lemma 3 to resolve such envies and obtain a new EFX allocation $Z$ such that $\phi(Z)=\phi(Y)>\phi(X)$.
@@
        $\mathbf{T=Y_{a_{i}}}$ for some $\mathbf{a_{i}\in\mathbb{A}}$: Then allocate $T$ to $c_{1}$ and $S^{\prime}$ to $a_{i}$. We then resolve internal strong envies within the groups using Lemma 3, yielding a new EFX allocation $Z$ such that $\phi(Z)=\phi(Y)>\phi(X)$.
@@
        $\mathbf{T=Y_{b_{j}}}$ for some $\mathbf{b_{j}\in\mathbb{B}}$ In this case, we ask agent $b_{j}$ to select her most preferred bundle from the agents in $\mathbb{A}$, say from agent $a_{i}$. We then allocate $Y_{a_{i}}$ to $b_{j}$, $Y_{b_{j}}$ to $c_{1}$, and $S^{\prime}$ to agent $a_{i}$. We know that agent $b_{1}$ does not strongly envy any agent from $\mathbb{A}\cup\mathbb{C}$. Thus, we swap the bundles of $b_{j}$ and $b_{1}$, so that now agent $b_{j}$ does not strongly envy anyone from $\mathbb{A}\cup\mathbb{C}$. Consequently, any strong envy towards agents in $\mathbb{A}\cup\mathbb{C}$, if it exists, must now come from agent $b_{1}$.
@@
      Agent $b_{1}$ $g$-champions $a_{1}$: Then, $\langle(a_{1},b_{i}),(b_{1},a_{1})\rangle$ forms a $\phi$-improving cycles. Therefore, using Lemma 3, we get a new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
@@
        Agent $b_{1}$ envies $c_{1}$: Here $\langle(a_{1},b_{i}),(b_{1},c_{1}),(c_{1},a_{1})\rangle$ is a $\phi$-improving cycle. Therefore, using Lemma 3, we get a new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
@@
          As $T_{c_{1}}^{a_{1}}\cup g$ is a minimally envied subset, no one strongly envies agent $a_{1}$. Furthermore, $Z_{a_{1}}>_{a}X_{a_{1}}>_{a}$, therefore agent $a_{1}$ is strictly better off compared to $X$ and $a_{1}$ does not strongly envy $c_{1}$. Therefore, we have an EFX allocation $Z$ such that $\phi(Z)>\phi(X)$. ∎
@@ 5.1.1 Overview of Algorithm 1
In this section, we give an overview of the algorithm. Please refer to Algorithm 1 the pseudo-code.
High-level idea: Recall that $a_{1}$ does not envy $b_{1}$ but $a_{1}$ $g$-champions $b_{1}$ with respect to the EFX allocation $X$. So the bundle $T_{b_{1}}^{a_{1}}\cup g$ is such that $a_{1}$ does not strongly envy anyone after receiving $T_{b_{1}}^{a_{1}}\cup g$ and in fact, this hold for any $a_{i}$, with respect to $X$. Further, $T_{b_{1}}^{a_{1}}\cup g>_{a}X_{a_{1}}$. Also, no one strongly envies $T_{b_{1}}^{a_{1}}\cup g$ with respect to $X$. If $T_{b_{1}}^{a_{1}}\cup g$ can be allotted to $a_{1}$, and a suitable replacement bundle for $b_{1}$ can be found, then this will result in an EFX allocation with improved value of $\phi$. This is roughly the goal of Algorithm 1. Temporarily, we allot $T_{b_{1}}^{a_{1}}\cup g$ to $b_{1}$. Clearly, the strong envy can only be from $b_{1}$ to other agents. This invariant, called Invariant $1$, is maintained throughout Algorithm 1.
@@ 2. Case 2
    $\mathbf{P\subseteq X_{c_{i}}}$ for some $\mathbf{i}$ and $\mathbf{Q\in\{X_{a_{j}}\mid a_{j}\in\mathbb{A}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}}$: See Figure 3. If $Q=X_{a_{j}}$ for some $j$, a similar improvement in $\phi$ can be obtained by giving $T_{b_{1}}^{a_{1}}\cup g$ to $a_{j}$, $P$ to $b_{1}$, and $Q$ to $c_{i}$. Note that this might need resolution of intra-group strong envies, which can be done by Lemma 4, and also sorting the bundles internally within each group so that the leading agents retain minimum valued bundles in their respective groups. A similar $\phi$ improvement can be made if $Q=T_{b_{1}}^{a_{1}}\cup g$ by using Lemma 8.
@@ 2. Case 3b
      $\mathbf{Q=P=X_{c_{i}}}$: Then, agents $b_{1}$ and $c_{i}$ compete for $P$ with $O_{b}$ and $O_{c}$ as their respective comparison sets. Each of $b_{1}$ and $c_{i}$ picks the minimum cardinality subsets, say $P_{b}$ and $P_{c}$ of $P$ respectively, such that, given this subset, they will have no envy towards any bundle in their respective comparison sets. We call $c_{i}$ to be the winner if $|P_{c}|\leq|P_{b}|$, and we call $P_{c}$ as the winning subset.
@@ 3. Case 3c
The correctness of Case 1 and Case 2 above is proved respectively in Part 1 and Part 2 of Lemma 10 below.

### v3

@@ 4 Three Sources in $E_{X}$
As the leading agents hold the minimum valued bundle in their respective groups, they envy all the agents in their own group and are never envied by anyone in their own group. We consider the case when no leading agent is envied by anyone. That is, $a_{1},b_{1}$ and $c_{1}$ are the sources in $E_{X}$.
@@
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If the leading agents are sources in the envy graph $E_{X}$, then there exists an EFX allocation $Y$ such that $Y\succ X$. Furthermore, $Y(L)\succ X(L)$ and $Y(\mathcal{N}\setminus{L})=X(\mathcal{N}\setminus{L})$.
@@
In this case, we only update the bundles of the leading agents, while the remaining agents retain their old bundles. After the update, each leading agent prefers their new bundle over their old ones. Additionally, the leading agents do not strongly envy the new bundles of other leading agents, even when compared to their own old bundles. Thus, the non-leading agents in a group (who keep their old bundles) do not strongly envy any agent in other groups. As a result, the only strong envies that remain are within the groups, which are resolved using Lemma 3. Therefore, the new allocation we obtain is an EFX allocation. Furthermore, the new bundles satisfy $Y(L)\succ X(L)$ and $Y(\mathcal{N}\setminus L)=X(\mathcal{N}\setminus L)$.
Proof of Lemma 6.
From Proposition 4, we know that the $g$-champion graph $M_{X}$ must contain a cycle of length at most three, involving agents $a_{1},b_{1}$, and $c_{1}$. If any of the leading agents self $g$-champions, using Corollary 1, we get a Pareto-dominating EFX allocation $Y$. If the $g$-champion cycle is of length $2$, then by applying Lemma 5, we get our desired allocation. Therefore, assume that $M_{X}$ has a cycle of length $3$. Without loss of generality, assume that $(a_{1},c_{1})$,$(c_{1},b_{1})$ and $(b_{1},a_{1})$ are $g$-champion edges in $M_{X}$. Since these agents do not envy each other, the $g$-champion edges break the bundle into top and bottom halves as shown below.
@@ 1. Case 1
    We allocate $T_{c}\cup B_{b}$ to agent $a_{1}$, $T_{a}\cup B_{c}$ to agent $b_{1}$, and $T_{b}\cup B_{a}$ to agent $c_{1}$. These bundles correspond to the most preferred top and bottom halves for each agent (refer to Conditions 4, 5, and 6). Therefore, all three agents strictly improve. Furthermore, according to agent $a_{1}$, we have $T_{c}\cup B_{b}>_{a}T_{a}\cup B_{a}=X_{a_{1}}>_{a}T_{a}\cup B_{c}=Y_{b_{1}}$. That is, $a_{1}$ does not envy $b_{1}$ even with respect to her old bundle $X_{a_{1}}$. Therefore, no agent in $\mathbb{A}$ envy $b_{1}$. Similarly, we have $X_{a_{1}}>_{a}T_{b}\cup B_{a}=Y_{c_{1}}$. That is, no agent from $\mathbb{A}$ envies $c_{1}$. Similarly we can show that no agents from either $\mathbb{B}$ or $\mathbb{C}$ envy any of $a_{1},b_{1}$ or $c_{1}$. If any strong envy remains, it must be confined to agents within their respective groups. Hence, we apply Lemma 3 to resolve these envies.
@@ 2. Case 2
    None of the three predicates holds: Thus $B_{a}>_{a}B_{b}$, $B_{b}>_{b}B_{c}$, and $B_{c}>_{c}B_{a}$. We then cyclically shift the top halves of these three bundles, allocating $T_{c}\cup B_{a}$ to agent $a_{1}$, $T_{a}\cup B_{b}$ to agent $b_{1}$, and $T_{b}\cup B_{c}$ to agent $c_{1}$.
    Now, since $T_{c}>_{a}T_{a}$ we know that $Y_{a_{1}}=T_{c}\cup B_{a}>_{a}T_{a}\cup B_{a}=X_{a_{1}}$. Therefore, agent $a_{1}$ strictly improved in $Y$. Similarly, we can show that agent $b_{1}$ and agent $c_{1}$ also strictly improves in $Y$.
    If there is any strong envy in $Y$, it must be directed towards or from the leading agents, as rest of the bundles remain unchanged. From case assumption, we have $B_{a}>_{a}B_{b}$. Therefore $X_{a_{1}}=T_{a}\cup B_{a}>_{a}T_{a}\cup B_{b}=Y_{b_{1}}$. As for each agent $a_{i}\in\mathbb{A}$, $Y_{a_{i}}\geq_{a}X_{a_{1}}>_{a}Y_{b_{1}}$, none of the agents in $\mathbb{A}$ envy $b_{1}$.
    Now we show that no agent in $\mathbb{A}$ envies $c_{1}$. Agent $a_{1}$ did not envy $c_{1}$ in $X$. Therefore, $X_{a_{1}}>_{a}T_{c}\cup B_{c}$. As $T_{c}>_{a}T_{b}$ (Condition 1), we have $X_{a_{1}}>_{a}T_{c}\cup B_{c}>_{a}T_{b}\cup B_{c}=Y_{c_{1}}$. Therefore, $a_{1}$ does not envy $c_{1}$. Furthermore, as $\forall a_{i}\in A,\ Y_{a_{i}}\geq_{a}X_{a_{i}}$, no agent in $\mathbb{A}$ envies $c_{1}$.
    With a symmetric argument, we can show that no agents in $\mathbb{B}$ or $\mathbb{C}$ envies any of the leading agents. Therefore, $Y$ is the required EFX allocation.
@@ 4. Case 4
        $\mathbf{S>_{c}X_{c_{1}}}$: Then, we allocate the bundles as follows. Agent $a_{1}$ gets $T_{c}\cup B_{a}$, $b_{1}$ retains $T_{b}\cup B_{b}$ and $c_{1}$ gets $S$. From Condition 4 we have $T_{c}\cup B_{a}>_{a}T_{a}\cup B_{a}$ and from this case assumption we know that $S>_{c}X_{c_{1}}$. Therefore agents $a_{1}$ and $c_{1}$ receive strict improvements while agent $b_{1}$ retains her old bundle.
@@
Let $X$ be a partial EFX allocation with at least one unallocated good $g$. If the leading agents in $L=\{a_{1},b_{1},c_{1}\}$ are all the sources in the envy graph $E_{X}$, then there exists an EFX allocation $Y$, possibly partial, such that at least one of the following holds:
@@ 5 Agent $a_{1}$ Envies someone in $\mathbb{B}\cup\mathbb{C}$
If, however, $b_{1}$ $g$-champions $a_{1}$, then $\langle(a_{1},b_{1}),(b_{1},a_{1})\rangle$ is a pseudo-cycle, thus we are done. The only remaining possibility is that $c_{1}$ $g$-champions $a_{1}$. We now consider cases based on who $g$-champions $c_{1}$.
@@ 3. Case 3
      Agent $c_{1}$ envies some $a_{i}\in\mathbb{A}\setminus\{a_{1}\}$: Then, we exchange the bundles of agents $a_{i}$ and $c_{1}$. Note that there may not be envy from $a_{i}$ to $c_{1}$, But since $a_{1}$ envies $c_{1}$, the bundle of $c_{1}$ is better than bundle of $a_{1}$ according to $v_{a}$. Now, all the strong envies are within the groups. Therefore, we apply Corollary 2 to resolve such envies and obtain a new EFX allocation $Z$ such that $\phi(Z)=\phi(Y)>\phi(X)$.
@@
        $\mathbf{T=Y_{a_{i}}}$ for some $\mathbf{a_{i}\in\mathbb{A}}$: Then allocate $T$ to $c_{1}$ and $S^{\prime}$ to $a_{i}$. We then resolve internal strong envies within the groups using Corollary 2, yielding a new EFX allocation $Z$ such that $\phi(Z)=\phi(Y)>\phi(X)$.
@@
        $\mathbf{T=Y_{b_{j}}}$ for some $\mathbf{b_{j}\in\mathbb{B}}$: In this case, we ask agent $b_{j}$ to select her most preferred bundle from the agents in $\mathbb{A}$, say from agent $a_{i}$. We then allocate $Y_{a_{i}}$ to $b_{j}$, $Y_{b_{j}}$ to $c_{1}$, and $S^{\prime}$ to agent $a_{i}$. We know that agent $b_{1}$ does not strongly envy any agent from $\mathbb{A}\cup\mathbb{C}$. Thus, we swap the bundles of $b_{j}$ and $b_{1}$, so that now agent $b_{j}$ does not strongly envy anyone from $\mathbb{A}\cup\mathbb{C}$. Consequently, any strong envy towards agents in $\mathbb{A}\cup\mathbb{C}$, if it exists, must now come from agent $b_{1}$.
@@
      Agent $b_{1}$ $g$-champions $a_{1}$: Then, $\langle(a_{1},b_{i}),(b_{1},a_{1})\rangle$ forms a pseudo-cycle. Therefore, using Lemma 4, we get a new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
@@
        Agent $b_{1}$ envies $c_{1}$: Here $\langle(a_{1},b_{i}),(b_{1},c_{1}),(c_{1},a_{1})\rangle$ is a pseudo-cycle. Therefore, using Lemma 4, we get a new EFX allocation $Y$ such that $\phi(Y)>\phi(X)$.
@@
          As $T_{c_{1}}^{a_{1}}\cup g$ is a minimally envied subset, no one strongly envies agent $a_{1}$. Furthermore, $Z_{a_{1}}>_{a}X_{a_{1}}>_{a}Z_{c_{1}}$, therefore agent $a_{1}$ is strictly better off compared to $X$ and $a_{1}$ does not strongly envy $c_{1}$. Therefore, we have an EFX allocation $Z$ such that $\phi(Z)>\phi(X)$. ∎
@@ 6.1.1 Overview of Algorithm 1
In this section, we give an overview of the algorithm. Please refer to Algorithm 1 for the pseudocode.
High-level idea: Recall that $a_{1}$ does not envy $b_{1}$ but $a_{1}$ $g$-champions $b_{1}$ with respect to the EFX allocation $X$. So the bundle $T_{b_{1}}^{a_{1}}\cup g$ is such that $a_{1}$ does not strongly envy anyone after receiving $T_{b_{1}}^{a_{1}}\cup g$ and in fact, this hold for any $a_{i}$, with respect to $X$. Further, $T_{b_{1}}^{a_{1}}\cup g>_{a}X_{a_{1}}$. Also, no one strongly envies $T_{b_{1}}^{a_{1}}\cup g$ with respect to $X$. If $T_{b_{1}}^{a_{1}}\cup g$ can be allotted to $a_{1}$, and a suitable replacement bundle for $b_{1}$ can be found, then this will result in an EFX allocation with improved value of $\phi$. This is roughly the goal of Algorithm 1. Temporarily, we allot $T_{b_{1}}^{a_{1}}\cup g$ to $b_{1}$. Clearly, the strong envy can only be from $b_{1}$ to other agents. This invariant, called Invariant $1$, holds throughout Algorithm 1.
@@ 2. Case 2
    $\mathbf{P\subseteq X_{c_{i}}}$ for some $\mathbf{i}$ and $\mathbf{Q\in\{X_{a_{j}}\mid a_{j}\in\mathbb{A}\}\cup\{T_{b_{1}}^{a_{1}}\cup g\}}$: See Figure 3. If $Q=X_{a_{j}}$ for some $j$, a similar improvement in $\phi$ can be obtained by giving $T_{b_{1}}^{a_{1}}\cup g$ to $a_{j}$, $P$ to $b_{1}$, and $Q$ to $c_{i}$. Note that this might need resolution of intra-group strong envies, using Lemma 3, and also sorting the bundles internally within each group so that the leading agents retain minimum valued bundles in their respective groups. A similar $\phi$ improvement can be made if $Q=T_{b_{1}}^{a_{1}}\cup g$ by using Lemma 8.
@@ 2. Case 3b
      $\mathbf{Q=P=X_{c_{i}}}$: Then, agents $b_{1}$ and $c_{i}$ compete for $P$ with $O_{b}$ and $O_{c}$ as their respective comparison sets. Each of $b_{1}$ and $c_{i}$ picks the minimum cardinality subsets, say $P_{b}$ and $P_{c}$ of $P$ respectively, such that, given this subset, they will have no envy towards any bundle in their respective comparison sets. We call $c_{i}$ to be the winner if $|P_{c}|\leq|P_{b}|$, and we call $P_{c}$ as the winning subset. For a formal definition of the competition framework, please refer to Definition 13.
@@ 3. Case 3c
The correctness of Case 1 and Case 2 above is proved respectively in Parts 1 and 2 of Lemma 10.
Definition 13.
Let $a,b\in\mathcal{N}$ be two agents, and let $O_{a}$ and $O_{b}$ represent the compare sets of agents $a$ and $b$ respectively, where $O_{a},O_{b}\subseteq 2^{\mathcal{M}}$ are collections of bundles of goods. A competition occurs when both agents prefer a given bundle $S\subseteq\mathcal{M}$ over all bundles in their respective compare sets. Formally, agents $a$ and $b$ compete for $S$ if:
$S>_{a}T\quad\forall T\in O_{a}\quad\text{and}\quad S>_{b}T\quad\forall T\in O_{b}.$
For each agent $i\in\{a,b\}$, let $P_{i}\subseteq S$ denote the smallest subset of $S$ that agent $i$ strictly prefers over every bundle in their compare set $O_{i}$. The agent with the smaller preferred subset is called the winner of the competition. Specifically, agent $a$ is the winner if $|P_{a}|\leq|P_{b}|$, and the loser otherwise.